in an oblique triangle, find the side b if c =12, A = 50 degree 12', and B = 72 degree 58'
Answer by stanbon(75887)

in an oblique triangle, find the side b if c =12, A = 50 degree 12', and B = 72 degree 58'
--------------------------
If A = 50 deg 12'
and B = 72 deg 58'
then C = 180 - (A+B)
= 180 - (123 10')
---
= 56 50'
----
Law of Sines:
b /sin(B) = c/sin(C)
b = sin(B)[c/sin(C)
b = sin(73.97)[12/sin(56.83)
b = 13.78
====================================
Cheers,
Stan H.